If one number exceeds another number by 13 and the larger number is 3/2 times the smaller number,then the smaller number is?
A)13 b)26 c)31 d)39

Answer by robertb(5830)

y = x + 13
y = 3x/2
==> 3x/2 = x + 13
==> x/2 = 13
==> x = 26
Letter (b).